# Spatiotemporal patterns in the active cyclic Potts model

### v1

@@ VI.2.1 Numerical Analysis
We first investigated the dynamics in cyclic-symmetry conditions, with $h_{01}=h_{12}=h_{20}=0.3$, in which case the three states are equivalent and the dynamics promotes the sequence $s=0\to 1\to 2\to 0$. We show in Fig. 14(b) that a biphasic $s=1,2$ band in a $s=0$ background forms a travelling wave, which behaves as a perfect soliton. The biphasic band is initialized with $(\rho_{1},\rho_{2})$ equals $(0,0)$ for the $s=0$ (gray) phase, $(0.8,0)$ for the $s=1$ (blue) phase and $(0,0.8)$ for the $s=2$ (red) phase. The stationary densities $(\rho_{1},\rho_{2})$ of the three phases are indicated as the green points in Fig. 14(a), matching the minima of the free energy $f$. Such travelling waves stochastically appear also in two dimensions, as can be seen in Fig. 3(a) and Movie S1. They also produce the spiralling dynamics around the contact point of three different domains (see Figs. 3(a) and 10(d)).
Biphasic bands travelling in opposite directions annhilate each other, as shown in Fig. 14(c), even when their sizes are different, as a natural consequence of the $s=0\to 1\to 2\to 0$ transformation. Such events can also be seen in two dimensions (see Fig. 10(d) and Movie S1).
In the case of asymmetric flipping energies, we find that three interfaces $s=0\to 1$, $s=1\to 2$ and $s=2\to 0$ have different velocities. Depending on the relative asymmetries, this can lead to the widening of a biphasic band, as shown in Fig. 15(a), or its disappearance (see Fig. 15(b)).

### v2

@@ VI.2.1 Numerical Analysis
We first investigated the dynamics in cyclic-symmetry conditions, with $h_{01}=h_{12}=h_{20}=0.3$, in which case the three states are equivalent and the dynamics promotes the sequence $s=0\to 1\to 2\to 0$. A biphasic $s=1,2$ band in a $s=0$ background forms a travelling wave, which behaves as a perfect soliton Noguchi et al. (2024). The stationary densities $(\rho_{1},\rho_{2})$ of the three phases are indicated as the green points in Fig. 15(a), matching the minima of the free energy $f$. Such travelling waves stochastically appear also in two dimensions, as can be seen in Fig. 3 and Movie S1. They also produce the spiralling dynamics around the contact point of three different domains (see Figs. 3 and 11(d)).
Biphasic bands travelling in opposite directions annhilate each other, as shown in Fig. 15(b), even when their sizes are different, as a natural consequence of the $s=0\to 1\to 2\to 0$ transformation. Such events can also be seen in two dimensions (see Fig. 11(d) and Movie S1).
In the case of asymmetric flipping energies, we find that three interfaces $s=0\to 1$, $s=1\to 2$ and $s=2\to 0$ have different velocities. Depending on the relative asymmetries, this can lead to the widening of a biphasic band or its disappearance (see Fig. 15(c)).
@@ VII Summary
Very recently, Manacorda and Fodor proposed a different type of thee-state lattice model Manacorda and Fodor . Their model allows the overlap of particles in each site and particles interact only with the other particles in the same site. The correlation of the neighbor sites is caused only by particle diffusion. They reported spiral wave and oscillatory modes but not homogeneous cycling mode. However, in the parameter range they call the arrest phase (one state covers the entire lattice), the homogeneous cycling might appear in longer simulations of their model.